Adam makes specialized garden figurines in a small shop on his property, and his monthly total sales revenue is $760 when he charges $20 for each figurine. One month, he tried lowering his price to $19, and his total sales revenue that month was $760. On the basis of these data, what is the price elasticity of demand for Adam’s product? Round your answer to 2 decimal places.
Answer #1

Q = Total sales revenue / Price

Q when price is 20 = 760 / 20 = 38

Q when price is 19 = 760 / 19 = 40

Elasticity of demand = [(q2-q1) /((q2+q1)/2)] / [(p2-p1) / ((p2+p1)/2)]

= [(40-38) /((40+38)/2)] / [(19-20) / ((19+20)/2)]

= [2 / 39] / [(-1) / 19.5]

= -1

Absslute value of elasicity = 1 (unit elastic)